OpenStudy (anonymous):

A rocket is shot upward from a 20 foot tall tower at 564 feet/sec. What is it's maximum height, when does it hit the ground and what is it's velocity at 10 seconds?

OpenStudy (cwrw238):

maxm. height is when final velocity = 0 v^2 = u^2 + 2 a s 0 = 564^2 - 2 * 32 s s = 564^2 / 64 = 4970. 25 ft above the tower time it hits ground use the equation s = ut + 0.5at^2 s = displacement = - 20 ft, a = -g = -32, t = ? -20 = 564t - 0.5*-32 t^2 16t^2 + 564t + 20 = 0 solve for t to give time in seconds it hits the ground

OpenStudy (cwrw238):

error - equation is 16t^2 - 564t - 20 = 0 t = 35.3 secs

OpenStudy (cwrw238):

velocity at 10 secs: t=10, u = 564, v = ?, a = -32 v = u + a t v = 564 - 32*10 v = 564 - 320 = 240 ft / sec = velocity after 10 seconds